OpenStudy (anonymous):

Find the equation of the line passing through the point (-6,13) which is parallel tot he line 8x+3y-14=0

OpenStudy (anonymous):

\[3y=-8x+14\] \[y=-\frac{8}{3}x+\frac{14}{3}\] \[m=-\frac{8}{3}x\] \[y-y_1=m(x-x_1)\] \[y-13=-\frac{8}{3}(x+6)\] \[3y-39=-8(x+6)\] \[3y-39=-8x-6\] \[3y+8x+9=0\]

ganeshie8 (ganeshie8):

its perfect !!

OpenStudy (anonymous):

Thanks for your response!